A stone of mass 2kg is released from a catapult whose rubber has been stretched through 40mm. If the force constant of the rubber is 200N/m. Calculate the velocity with which the stone leaves the catapult?
1
Expert's answer
2020-07-15T09:35:42-0400

"\frac{kx^2}{2}=\frac{mv^2}{2}\to v=\sqrt{\frac{kx^2}{m}}=x\sqrt{\frac{k}{m}}=0.04\sqrt{\frac{200}{2}}=0.4m/s"